### v1

# Importance-based Token Merging for Diffusion Models

Abstract
Diffusion models excel at high-quality image and video generation. However, a major drawback is their high latency. A simple yet powerful way to speed them up is by merging similar tokens for faster computation, though this can result in some quality loss. In this paper, we demonstrate that preserving important tokens during merging significantly improves sample quality. Notably, the importance of each token can be reliably determined using the classifier-free guidance magnitude, as this measure is strongly correlated with the conditioning input and corresponds to output fidelity. Since classifier-free guidance incurs no additional computational cost or requires extra modules, our method can be easily integrated into most diffusion-based frameworks. Experiments show that our approach significantly outperforms the baseline across various applications, including text-to-image synthesis, multi-view image generation, and video generation.
1 Introduction
Diffusion models [26, 89] excel in image and video generation [78, 79, 2, 17]. The rise of powerful diffusion models such as DALL-E [74], Stable Diffusion (SD) [78], or Imagen [79] has dramatically changed the landscape of generative AI. At their core, these models operate by iteratively denoising images or videos through multiple passes of a backbone network, processing a substantial number of tokens, which makes them particularly computationally intensive. For example, it could take up to 6 seconds to generate a single image with the Stable Diffusion model. A simple but powerful approach to speed up is token merging [4, 3] (ToMe), which merges similar or redundant tokens before executing computationally heavy operations, such as attention, making it highly effective for generating high-resolution images or videos.
ToMe frameworks typically divide tokens into two sets: source (src) and destination (dst) where most similar tokens from the src set are merged into the dst set. Essentially, tokens in the dst set function as anchors for grouping related tokens and also serve as the primary regions where computation will take place. Thus, token partitioning is particularly crucial to minimize performance drop and avoid image artifacts. To this end, ToMeSD [3] proposes to select a destination token randomly within each $2\times 2$ region, shown to be better than both fully random sampling and other fixed-pattern methods. However, this selection scheme is still random in nature and does not prioritize tokens based on relevance, which potentially misallocates resources to non-essential areas, as can be seen in Fig. 1 (a).
Wise old man reading a book by candlelightRefer to caption (a) ToMeSD [3] (b) Ours
Figure 1: Importance-based Token Merging. Our method prioritizes important tokens during token merging, resulting in images with greater details in essential areas compared to ToMeSD [3]. In the second row, we show the regions (in white) where computation (e.g., attention) will take place after token merging.
To address these limitations, we propose a novel token selection method that focuses computational resources on areas of high relevance to the prompt. Unlike the random selection used in ToMeSD [3], our approach leverages classifier-free guidance [25] as a targeted mechanism for token selection. While classifier-free guidance has been typically used as a global adjustment to emphasize prompt-relevant features, here we show that we can use it at the token level to dynamically highlight tokens carrying more essential information. By prioritizing token selection based on this prompt relevance, we ensure that high-information regions receive focused attention, preserving critical details and enhancing image quality. Further, classifier-free guidance incurs no additional computational cost and requires no modification to the architecture, thus, it can be readily used for any conditional diffusion models.
More specifically, our method assigns each token a relevance score based on its magnitude in the classifier-free guidance map, indicating its importance to the prompt. We then construct a pool of important tokens and only select destination tokens from this pool. Unlike simply picking the top-k important tokens, which can lead to redundancy and a suboptimal allocation of computing resources, our approach ensures a balance between relevance and diversity, focusing on prompt-aligned areas while maintaining efficiency. We apply our token merging strategy across three key applications: text-to-image generation, text-to-video generation, and multi-view image generation. In all cases, our method can be integrated seamlessly without requiring any modifications. Compared to the baseline method, our approach consistently demonstrates superior performance, delivering results with higher fidelity and significantly improved image details across all tested scenarios.
Our contributions are as follows:
  • We propose using classifier-free guidance magnitude as a measure of token importance. This can be obtained efficiently and robustly without any modification to diffusion models.
  • We design a novel token-merging strategy based on token importance to improve the quality of generated samples. We introduce an adaptive important token pool to enhance resource allocation.
  • Our method achieves state-of-the-art performance across various diffusion model tasks, including text-to-image synthesis, multi-view generation, and video generation.
2 Related Work
2.1 Diffusion Models
Diffusion models [87, 12, 26, 89] are a class of generative models that iteratively transform random noise into complex data structures, such as images, by gradually reversing a diffusion process. In these models, data is progressively corrupted by adding noise, and the model learns to recover the original data through iterative denoising. This approach has demonstrated impressive results in generating high-quality, realistic images [78, 79, 70, 6, 74, 66, 16, 64, 104, 1], videos [83, 2, 17, 113, 38, 5], 3D content [72, 51, 56, 82, 41, 59, 67], and audio [36, 27].
In diffusion models, classifier-free guidance (CFG) [25] is a technique that enhances fidelity and detail in generation by guiding the model toward conditioned inputs without the need for an external classifier. Recent work [109, 97] suggests a connection between CFG and saliency. Our study advances this by identifying CFG as an indicator of token importance, enabling more effective token merging.
To reduce the cost of diffusion inferences, various techniques have been proposed. Some approaches require retraining, including better model structures [78, 71, 34], model pruning [14], model compression [112, 105, 46], step distillation [80, 65, 81, 52, 18], and consistency regularization [90, 60]. Others avoid retraining, such as improved sampling to reduce inference steps [89, 50, 57, 58], caching [62, 100, 8, 111, 31, 85, 61], model quantization [44, 7, 22, 94, 11, 86], and token reduction [32, 3, 4]. In this work, we focus on improving token reduction. Notably, our method is compatible with other diffusion acceleration techniques, enabling combined use to further improve inference efficiency.
2.2 Token Reduction
Token reduction [20, 4, 92] decreases the number of tokens to process by pruning or merging them. It is widely applied in tasks such as classification [63, 4, 20, 48], segmentation [102, 33], detection [54], video understanding [10], and large language models [43, 30, 88]. Token pruning methods include removing tokens based on attention scores [15, 20, 48, 55, 99, 101, 103], using the Gumbel-Softmax trick for selective pruning [29, 37, 54, 75, 99], developing sampling methods [15, 106], integrating sparsity in the model [9, 42], and reinforcement learning-based methods [68]. For token merging [4, 19, 77, 102, 114, 91, 96, 39], approaches include bipartite soft matching [4, 3], K-Means [63], K-Medoids clustering [63], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) [107]. Additionally, soft merging methods assign tokens to multiple clusters before merging them [19, 77, 114].
Recent studies have applied token reduction to diffusion model inference [3, 35, 21, 93, 45, 32, 110, 47, 95, 84]. ToMeSD [3] introduced token merging to Stable Diffusion [78] with a training-free method. It selects destination tokens and uses bipartite soft matching to merge redundant tokens. The reduced token set is processed by operations like attention, and then copied back to merged token locations to ensure the final token count matches the input. While ToMeSD performs well, its selection of destination tokens is not optimal, i.e., random across the image. In our work, we propose that selecting destination tokens based on their importance improves the fidelity and quality in generation, and this importance can be easily obtained via classifier-free guidance without additional cost. AT-EDM [93] uses self-attention maps to derive token importance, but this requires first performing full self-attention and increases peak memory usage. ATC [21] applies bottom-up hierarchical clustering for better token merging, but our experiment shows that this clustering is costly for generative tasks.
3 Preliminaries
Diffusion models.
Diffusion models [12, 26, 89] are generative models that produce data by simulating a denoising process. Starting with data $\mathbf{x}_{0}$, the forward diffusion process gradually adds Gaussian noise over $T$ timesteps, producing a sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$. This process is defined as:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\,\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\beta_{t}$ controls the noise schedule. The reverse process denoises the input by modeling $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ with learned parameters. Most models predict the noise added at each timestep, $\epsilon_{\theta}(\mathbf{x}_{t},t)$.
Classifier-free guidance (CFG) [25] modifies the noise prediction to improve sample control. It is designed to steer the predicted noise in a direction more aligned with the condition:
$\tilde{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)),$ (2)
where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$, $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction, and $w$ is the guidance weight.
Token Merging for Diffusion Models.
Our token merging method builds on ToMeSD [3], which is designed to speed up text-to-image diffusion models. To accelerate a model layer, such as attention, ToMeSD divides the input feature tokens into a source (src) set and a destination (dst) set. The dst set is selected randomly, i.e., by choosing one token from each $2\times 2$ region in the input feature map. Using bipartite soft matching, it then selects and merges the $r$ most similar tokens from the src set into the dst set, where $r$ represents the token merging ratio. The merging is done via averaging. After the model layer processes this reduced token set, the merged locations are filled with the corresponding processed dst tokens, maintaining the same output shape as the input.
4 Method
Figure 2: Overview. We propose an importance-based token merging method. The importance of each token is determined by the magnitude of classifier-free guidance. These scores, visualized with colors ranging from light to dark (indicating less to more important tokens), are used to construct a pool of important tokens. We randomly select a set of destination (dst) tokens from this pool and the remaining important tokens become source (src) tokens. Bipartite soft matching is then performed between the dst tokens and src tokens. src tokens without a suitable match are considered independent tokens (ind.). All other src tokens and unimportant tokens are merged with the destination tokens for subsequent computational steps.
We propose a novel token-merging method for diffusion models. Our approach uses classifier-free guidance to identify important tokens, allowing for a more efficient allocation of computing resources. Additionally, we create a dynamic pool of important tokens, with the pool size adapting to the token merging ratio. This adjustment enhances resource allocation and reduces redundancy when determining which tokens to merge. Fig. 2 provides an overview of our method.
Important Tokens.
Selecting which tokens to serve as anchors (destination tokens) for merging is critical for generating high-quality content. This is because these tokens correspond to the primary image regions where subsequent computations, such as attention, are applied. However, ToMeSD [3] currently relies on a random selection process based on a simple heuristic. It lacks well-established criteria for determining token importance. The primary guiding principle of ToMeSD is to prioritize spatial uniformity among the destination tokens [3, 4].
Fresh citrus and berriesRefer to caption
Figure 3: Importance Maps. We show token importance maps derived from classifier-free guidance across diffusion inference timesteps. The generated image from the diffusion model is displayed on the left for reference.
In this paper, we highlight that classifier-free guidance [25] serves as a reliable indicator of token importance, providing a principled approach to identifying image regions with essential information. By leveraging this connection, we propose a method that prioritizes computing resources on important tokens, enabling more efficient and informed token merging even under constrained compute budgets. For each token, we calculate its importance as the absolute value of its classifier-free guidance score:
$\text{importance}=\left|\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)\right|,$ (3)
where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. In Fig. 3, we provide an example of the importance maps.
A car made out of vegetablesRefer to caption (a) Top-k (b) Ours
Figure 4: We compare our method with an approach that uses the top-k important tokens as destination tokens (dst) for token merging. The computation locations after token merging are illustrated as non-black pixels in the bottom-left windows. They include locations of dst tokens, which are shown in white, and independent tokens (some other tokens that lack a similar dst token for merging), which are shown in orange.
Figure 5: Qualitative comparison of text-to-image generation. The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with various token merging methods including ATC [21], ToMeSD [3], and our proposed method. Our approach consistently produces finer details with coherent structures. Note that ATC requires minutes to generate an image, whereas other methods, including ours, complete the task in seconds. The token merging ratio is 0.7. Please refer to the supplementary for detailed prompts. Best viewed with zoom-in for clarity.
Figure 6: Qualitative comparison of multi-view diffusion. We apply ToMeSD [3] and our token merging method to the multi-view diffusion model. We use Zero123++ [82] as the base model and a merging ratio of 0.6. Best viewed with zoom-in. Please refer to supplementary for input images.
Importance-based Token Merging.
With the token importance scores, a naive approach is to pick the top-k tokens as destination tokens and merge the rest tokens that are similar to them. However, this approach produces low-quality outputs due to merging inefficiency, as shown in Fig. 4 (a). More specifically, there are two particular issues about this:
  1. 1.
    Redundancy. The top-k tokens can be very similar - but all important tokens, leading to redundancy and less intra-variant among the selected destination tokens.
  2. 2.
    Unimportant Independent Tokens. In the standard token merging pipeline [3], there are “independent” tokens that don’t have similar destination tokens, which are processed independently without merging. In our case, background or irrelevant tokens can easily become independent tokens as there are no good matches for them among the important tokens, as shown in Fig. 4.
To this end, we propose to first create a pool of the most important tokens and ensure both destination tokens and independent tokens are drawn from this set. Destination tokens are randomly sampled from the pool, while independent tokens are selected as those in the important token pool that are most dissimilar to the chosen destination tokens. This simple method ensures that all subsequent computations after the merging step are performed on important tokens, leading to improved detail preservation, as demonstrated in Fig. 4 (b).
To determine the optimal pool size, we adapt it based on the token merging ratio $r$. Specifically, we set the pool size as $\mathbf{P}=(1-r)\cdot(1+p)$, where $p$ is a hyper-parameter of our method. With a constrained token processing budget, i.e., a high token merging ratio $r$, the pool size remains small, ensuring the selection of only the most critical tokens. Conversely, with a larger compute budget, the pool size increases, reducing the likelihood of selecting redundant tokens as destination tokens.
Token Merging in Diffusion Inference.
More specifically, at time-step $t$ of the diffusion inference, a diffusion model layer, e.g. a transformer layer, takes $\mathbf{N}$ tokens as input. The token merging ratio is $r$. Based on the token importance derived from the previous timestep’s classifier-free guidance, we select the top $\mathbf{K}=\mathbf{N}\cdot(1-r)\cdot(1+p)$ tokens as the important token pool, denoted as $\mathbf{A}$. From $\mathbf{A}$, we randomly pick $\mathbf{D}=\mathbf{N}\cdot k$ tokens to form the destination (dst) set. Here, $p,k$ are the two hyper-parameters of our method. The remaining tokens in $\mathbf{A}$ become the source (src) set.
Next, we perform bipartite soft matching by computing pairwise cosine similarities between src and dst tokens. Each src token is linked to its most similar dst. Then, in src set, we select the top $\mathbf{I}=\mathbf{N}\cdot(1-k-r)$ tokens with the smallest similarity to their closest dst tokens to serve as independent tokens. The remaining src tokens and unimportant tokens are merged into their corresponding dst tokens. After merging, the number of tokens reduces to $\mathbf{I}+\mathbf{D}=\mathbf{N}\cdot(1-r)$ tokens.
Preserving Structure in Early Time-steps.
Prior work [35] suggests that token pruning (directly dropping tokens) could help preserve structural details. Building on this observation, we found that incorporate token pruning during the early diffusion steps, followed by token merging in later steps, improves generation results. Furthermore, the low token variance in the early steps [93] reduces the effectiveness of classifier-free guidance in identifying important tokens. To mitigate this, we simply employ the ToMeSD strategy during these initial steps, e.g., the first 6 out of 50 steps, to select one token from each $2\times 2$ region of the feature map as a destination token.
5 Experiments
Our token merging strategy leverages classifier-free guidance for relevance-based token selection, ensuring an efficient and seamless integration without requiring any additional modules or modifications to diffusion models. To showcase its effectiveness, we evaluate our method across three diffusion generation tasks: text-to-image generation, image-conditioned multi-view diffusion, and text-to-video diffusion.
5.1 Experimental Settings
Text-to-image Generation.
We use Stable Diffusion 2 (SD) [78] as the base model, a text-to-image latent diffusion model with a U-Net architecture that generates 768x768 images. Our token merging method is compared to ToMeSD [3] and ATC [21]. Due to ATC’s slow inference (several minutes per image), we only display its visual results. For quantitative comparison, we follow previous studies [79, 104, 1, 74] and report FID [24] and CLIP scores [73, 23] for zero-shot image generation on the MS-COCO 2014 validation dataset [49], with 30K randomly sampled image-caption pairs.
Multi-view Diffusion.
We use Zero123++ v1.2 [82] as the base model. Zero123++ is an image-conditioned multi-view latent diffusion model that generates six novel views at a resolution of 320×320. We compare our token merging method to ToMeSD. Following the evaluation protocols of prior work [53, 56], we test on GSO dataset [13], which comprises 30 everyday objects, and compute PSNR, SSIM [98], and LPIPS metrics [108] to evaluate the similarity between the generated images and ground truth. Additionally, we include visual comparisons using in-the-wild images as input to further demonstrate the effectiveness of our method.
Video Diffusion.
We adopt AnimateDiff v3 [17] as the base model, which adds temporal attention layers to turn the text-to-image model, i.e. Stable Diffusion, into a video diffusion model. This model generates 16-frame videos at a resolution of 512×512. We compare our token merging method with ToMeSD and evaluate using VBench [28]. We report the semantic, quality, and total scores from VBench. The semantic score assesses alignment between the generated videos and the user prompt, focusing on entity types, attributes, and styles. The quality score evaluates the temporal consistency and visual quality of the generated videos.
Other Metrics.
We also report TFLOPs, latency, and GPU memory usage, with and without memory-efficient attention (xFormers [40]), on an NVIDIA A5000 GPU. Inference is performed using float16 precision. We estimate TFLOPs for a single diffusion sampling step. For latent diffusion models, the inference cost is measured exclusively in the latent space.
Implementation Details.
For text-to-image diffusion, we follow the approach used in ToMeSD [3] by merging tokens in the self-attention layers of the first and last blocks of Stable Diffusion. The diffusion process consists of 50 sampling steps, with token pruning applied during the first 6 steps, followed by token merging. The classifier-free guidance (CFG) scale is set to 7.5. For multi-view diffusion, we merge tokens in the self-attention layers of the first two and last two blocks of Zero123++ [82]. The process involves 50 sampling steps, with token pruning applied in the first 10 steps. The classifier-free guidance (CFG) scale is set to 4. For video diffusion, token merging is applied in the first and last blocks of AnimateDiff [17]. The sampling consists of 30 steps, with token pruning during the first 4 steps and a CFG scale of 7.5. To ensure fairness, these settings are consistently applied to both our method and the baselines. The hyper-parameter $p$, which determines the size of the important token pool, is set to 0.4, 0.6, and 0.8 for image, multi-view, and video generation task, respectively. The number of destination tokens ($k$) is consistent across all tasks and follows the setting used in ToMeSD, utilizing 25% of the total tokens.
5.2 Results
Figure 7: Qualitative comparison of video diffusion. We apply ToMeSD and our token merging method to the video diffusion model. For each generated video, we show three frames: the first on the left, the 8th at the top right, and the last 16th frame at the bottom right. We use AnimateDiff [17] as the base model and a merging ratio of 0.7. Best viewed with zoom-in. Please refer to supplementary for prompts.
$r$ FID $\downarrow$ CLIP $\uparrow$
ToMeSD Ours ToMeSD Ours
0 11.88 31.83
0.30 12.20 12.20 31.82 31.82
0.50 13.50 13.42 31.79 31.83
0.60 14.81 14.51 31.80 31.81
0.70 17.46 16.22 31.78 31.79
0.75 20.89 17.75 31.71 31.76
Table 1: Text-to-image generation. We apply ToMeSD [3] and our token merging method to Stable Diffusion [78] across various token merging ratios $r$.
In Tab. 1, we compare different token merging methods applied to Stable Diffusion 2 [78] for text-to-image generation. Our method consistently outperforms ToMeSD [3] in this task, especially at higher token merging ratios ($r$). For example, at $r=0.75$, our method achieves an FID of 17.75 versus ToMeSD’s 20.89. Agglomerative Token Clustering (ATC) [21] is not included due to its prohibitive computational cost, as it is CPU-bound and non-batched, making it impractical for large-scale evaluations. We provide qualitative comparisons in Fig. 5. Note that when the token merging ratio $r$ is small, such as 0.3, our important token pool becomes the whole token set, making it functionally equivalent to ToMeSD.
$r$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ToMe. Ours ToMe. Ours ToMe. Ours
0.40 14.80 14.82 0.775 0.777 0.260 0.259
0.60 14.71 14.83 0.782 0.783 0.272 0.269
0.70 14.18 14.80 0.787 0.785 0.302 0.274
0.75 13.12 14.58 0.789 0.784 0.349 0.283
Table 2: Multi-view diffusion. We apply ToMeSD [3] and our token merging method to Zero123++ [82] with different token merging ratios $r$.
In Tab. 2, we compare ToMeSD and our token merging method in the context of multi-view diffusion. Our importance-based token merging method consistently shows improved performance over ToMeSD, especially at higher merging ratios. At $r=0.75$, our method achieves a significant improvement in PSNR (14.58 vs. 13.12) and a much lower LPIPS (0.283 vs. 0.349), highlighting its ability to maintain high output quality even under aggressive token compression. Qualitative examples in Fig. 6 further validate this, showcasing finer geometrical and textual details in the objects generated by our method.
$r$ Semantic $\uparrow$ Quality $\uparrow$ Total $\uparrow$
ToMe. Ours ToMe. Ours ToMe. Ours
0.40 75.40 75.40 81.69 81.69 80.44 80.44
0.60 74.03 74.51 81.58 81.75 80.07 80.30
0.70 72.03 73.23 81.52 81.23 79.62 79.63
0.75 69.67 71.58 80.82 81.00 78.59 79.12
Table 3: Video diffusion (spatial). Token merging applies on only spatial attention layers of AnimateDiff [17] with various token merging ratios $r$. We compare our method and ToMeSD [3] with VBench scores [28].
Semantic $\uparrow$ Quality $\uparrow$ Total $\uparrow$
ToMeSD 72.71 81.35 79.62
Ours 73.52 81.69 80.06
Table 4: Video diffusion (spatial and temporal). Token merging applies on both spatial and temporal attention layers of AnimateDiff [17] with a spatial merging ratio of 0.7 and a temporal merging ratio of 0.2. We report VBench scores [28] of our method in comparison to ToMeSD [3].
A similar trend can be observed in Tab. 3 and Fig. 7, where we extend the comparison to video diffusion. Across these tests, our method consistently performs better than ToMeSD, both in numerical metrics and in visual quality, particularly in preserving object details. We observed that merging tokens in the temporal layers significantly reduces the generated dynamics. Nonetheless, we present results with token merging applied for both spatial and temporal layers in Tab. 4. In summary, our method outperforms ToMeSD across all three tasks, generating images and videos with better object details.
$r$ Latency (s) Memory (GB) TFLOPs
ToMeSD Ours
0 8.5 (5.3) 7.63 (3.16) 4.30
0.3 8.0 (5.0) 8.0 (5.0) 5.20 (3.16) 3.83
0.5 6.0 (4.7) 6.1 (4.8) 4.05 (3.16) 3.55
0.7 5.7 (4.5) 5.8 (4.5) 3.55 (3.16) 3.36
Table 5: Comparison of inference costs when applying our token merging method and ToMeSD [3] to Stable Diffusion 2 [78]. Numbers in parentheses indicate measurements when memory-efficient attention [40] is enabled.
Inference Costs.
We compare the inference costs of our method and ToMeSD when applied to Stable Diffusion 2 in Tab. 5. As can be seen, both methods show similar improvements in inference times, GPU usage, and TFLOPs. This demonstrates that our token merging strategy can enhance performance without incurring additional computational costs.
5.3 Ablation Study
$r$ Ours w/ top-k dst w/ global ind.
0.3 FID $\downarrow$ 12.20 12.56 12.22
CLIP $\uparrow$ 31.82 31.82 31.82
0.6 FID $\downarrow$ 14.51 14.56 14.73
CLIP $\uparrow$ 31.81 31.83 31.82
Table 6: Ablation studies on alternative design choices of our method for text-to-image generation. ’w/ top-k dst’ means top-k rather than random selection from the important token pool for destination tokens. ’w/ global ind.’ means independent tokens may also be outside the important token pool, instead of solely within it.
$r$ $p$ 0 0.2 0.4 0.8
0.3 FID $\downarrow$ 12.87 12.32 12.19 12.19
CLIP $\uparrow$ 31.83 31.83 31.82 31.82
0.7 FID $\downarrow$ 16.52 16.23 16.22 16.42
CLIP $\uparrow$ 31.78 31.78 31.79 31.79
Table 7: Choice of $p$. We show the results of our method for the text-to-image generation task with different values of $p$, which determines the important token pool size.
We conduct ablation studies on the text-to-image generation task to examine alternative design choices in our method. As shown in Tab. 6, simply using top-k important tokens as destination tokens leads to worse results. Furthermore, not choosing independent tokens explicitly from the important set (w/ global ind.), leads to a performance drop.
Choice of $p$.
As described in Sec. 4, the size of the important token pool is calculated as $(1+r)\cdot(1+p)$, where $r$ represents the token merging ratio. In Tab. 7, we present results for text-to-image generation using different values of $p$. Setting $p$ too low results in an overly small pool of important tokens, which increases redundancy when randomly selecting destination tokens from it. Conversely, setting $p$ too high includes too many unimportant tokens in the pool, negatively impacting performance.
6 Conclusions
In this paper, we propose an importance-based token merging method for diffusion models, which maintains generation quality while reducing inference latency. Leveraging classifier-free guidance as an indicator of token importance, our approach strategically allocates computational resources to regions of high relevance to the input condition, thereby enhancing the fidelity of the generated outputs. Our method achieves state-of-the-art performance across diverse tasks, including text-to-image synthesis, multi-view generation, and video generation, highlighting its effectiveness and versatility. Our approach reveals the surprising effect of applying classifier-free guidance at the token level to facilitate compression with minimal impact on quality.
Supplementary Material
In Appendix A, we present more qualitative comparisons, results on preserving output structure, and experiments with varying diffusion inference steps. In Appendix B, we provide more details about our experimental settings. In Appendix C, we discuss the limitations of our method. In Appendix D, we provide the prompts used to generate qualitative results. We also include a supplementary video for comparisons on text-to-video generation.
Appendix A Additional Results
Additional Qualitative Results.
We provide additional qualitative comparisons between our token merging method and ToMeSD [3]. Additional visual comparisons for text-to-image generation are shown in Fig. 8, and results for multi-view diffusion are presented in Fig. 9. In the supplementary video, we include comparisons on text-to-video generation, using AnimateDiff [17] as the base diffusion model and a merging ratio of 0.7. Furthermore, we provide visual comparisons between ToMeSD [3] and our method across various merging ratios for text-to-image generation in Fig. 10.
Preserving Structure in Early Timesteps.
In Tab. 8 and Fig. 11, we compare the results of ToMeSD [3] under two scenarios: (1) applying token pruning during the early diffusion steps followed by token merging, and (2) using token merging throughout all diffusion steps. The results show that token pruning in the early steps more effectively preserves the generation layout.
$r$ FID $\downarrow$ CLIP $\uparrow$
w/o pr. w/ pr. w/o pr. w/ pr.
0.10 11.75 11.72 31.81 31.82
0.30 12.16 12.20 31.82 31.82
0.50 13.49 13.50 31.79 31.79
0.60 14.81 14.81 31.79 31.80
0.70 17.51 17.46 31.76 31.78
0.75 21.05 20.89 31.69 31.71
Table 8: Ablation studies on token pruning in early diffusion inference steps. We compare the results of ToMeSD [3] with token pruning in early diffusion inference steps followed by token merging, versus using token merging for all steps. We evaluate using the text-to-image generation task with Stable Diffusion [78] as the base model across various token merging ratios $r$.
Number of Diffusion Inference Steps
In Tab. 9, we compare ToMeSD [3] and our method across different numbers of diffusion inference steps for text-to-image generation, demonstrating that our method consistently outperforms ToMeSD.
$T$ FID $\downarrow$ CLIP $\uparrow$
ToMeSD Ours ToMeSD Ours
20 18.57 17.03 31.77 31.80
30 17.82 16.51 31.78 31.80
50 17.46 16.22 31.78 31.79
Table 9: Number of Diffusion Inference Steps. We compare ToMeSD [3] and our token merging method for the text-to-image generation task with different diffusion inference steps, using Stable Diffusion [78] as the base model and a merging ratio of 0.7.
Figure 8: Additional comparison of text-to-image generation. The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with ToMeSD [3] and our token merging method. The token merging ratio is 0.7. Best viewed with zoom-in for clarity.
Figure 9: Additional qualitative comparison of multi-view diffusion. We apply ToMeSD [3] and our token merging method to the multi-view diffusion model. We use Zero123++ [82] as the base model and a merging ratio of 0.6. Best viewed with zoom-in.
Figure 10: We provide an additional comparison between ToMeSD [3] and our method when applied to Stable Diffusion [78] across various merging ratios $r$. For reference, the results of Stable Diffusion without token merging are shown on the left. Best viewed with zoom-in for clarity.
Figure 11: We compare the results of ToMeSD [3] with token pruning in early diffusion inference steps followed by token merging (w/ prune), versus using token merging for all steps (w/o prune). We use Stable Diffusion [78] as the base model and a merging ratio of 0.6.
Appendix B Additional Experimental Settings
Metrics.
We use the deepspeed [76] library to estimate TFLOPs, the clean-fid [69] library to calculate FID scores, and the openai/clip-vit-base-patch16 model from OpenAI-CLIP [73] to calculate CLIP scores.
Details on Multi-view Diffusion.
The base multi-view diffusion model used in our experiments is Zero123++ [82], which fine-tunes Stable Diffusion 2 [78] to generate six novel views from an input image. During denoising, the model appends the self-attention key and value matrices from the reference input image to the attention layers for conditioning. The novel view poses are defined by a fixed set of absolute elevation and relative azimuth angles. Specifically, the elevation and azimuth angles (in degrees) are set as follows: (30, 30), (-20, 90), (30, 150), (-20, 210), (30, 270), (-20, 330). We consistently use this sequence of novel views to present visual results in multi-view diffusion experiments.
Appendix C Discussion and Limitations
Our method demonstrates broad applicability across diffusion models. However, step-distilled diffusion models [65] do not explicitly compute classifier-free guidance, as this information is already distilled into the final model. An interesting direction for future research could involve refining the distillation process to enable the model to predict an additional output: a classifier-free guidance map, which could then be used for better token merging or other innovative applications.
Appendix D Prompts
We provide the prompts used to generate the qualitative results shown in the paper but not included in the figures.
Text prompts corresponding to the text-to-image generations in Fig. 5:
  • Colorful butterfly with wings fully spread
  • Elegant teacup with a delicate floral pattern
  • Royal castle with golden towers at sunrise
  • Young musician playing guitar on stage
Text prompts corresponding to the text-to-video generations in Fig. 7:
  • Tower
  • A jellyfish floating through the ocean, with bioluminescent tentacles
  • In a still frame, the ornate Victorian streetlamp stands solemnly, adorned with intricate ironwork and stained glass panels
In Fig. 12, we show image prompts corresponding to the image-conditioned multi-view generations in Fig. 6.
Figure 12: Image prompts for Fig. 6.
In Fig. 13, we show image prompts corresponding to the image-conditioned multi-view generations in Fig. 9.
Figure 13: Image prompts for Fig. 9.

### v2

@@
-Importance-based Token Merging for Diffusion Models
+Importance-Based Token Merging for Efficient Image and Video Generation
 Abstract
-Diffusion models excel at high-quality image and video generation.
-However, a major drawback is their high latency.
-A simple yet powerful way to speed them up is by merging similar tokens for faster computation, though this can result in some quality loss.
-In this paper, we demonstrate that preserving important tokens during merging significantly improves sample quality.
-Notably, the importance of each token can be reliably determined using the classifier-free guidance magnitude, as this measure is strongly correlated with the conditioning input and corresponds to output fidelity.
-Since classifier-free guidance incurs no additional computational cost or requires extra modules, our method can be easily integrated into most diffusion-based frameworks.
-Experiments show that our approach significantly outperforms the baseline across various applications, including text-to-image synthesis, multi-view image generation, and video generation.
+Token merging can effectively accelerate various vision systems by processing groups of similar tokens only once and sharing the results across them.
+However, existing token grouping methods are often ad hoc and random, disregarding the actual content of the samples.
+We show that preserving high-information tokens during merging—those essential for semantic fidelity and structural details—significantly improves sample quality, producing finer details and more coherent, realistic generations.
+Despite being simple and intuitive, this approach remains underexplored.
+To do so, we propose an importance-based token merging method that prioritizes the most critical tokens in computational resource allocation, leveraging readily available importance scores, such as those from classifier-free guidance in diffusion models.
+Experiments show that our approach significantly outperforms baseline methods across multiple applications, including text-to-image synthesis, multi-view image generation, and video generation with various model architectures such as Stable Diffusion, Zero123++, AnimateDiff, or PixArt-$\alpha$.
 1 Introduction
-Diffusion models [26, 89] excel in image and video generation [78, 79, 2, 17].
-The rise of powerful diffusion models such as DALL-E [74], Stable Diffusion (SD) [78], or Imagen [79] has dramatically changed the landscape of generative AI.
-At their core, these models operate by iteratively denoising images or videos through multiple passes of a backbone network, processing a substantial number of tokens, which makes them particularly computationally intensive.
-For example, it could take up to 6 seconds to generate a single image with the Stable Diffusion model.
-A simple but powerful approach to speed up is token merging [4, 3] (ToMe), which merges similar or redundant tokens before executing computationally heavy operations, such as attention, making it highly effective for generating high-resolution images or videos.
-ToMe frameworks typically divide tokens into two sets: source (src) and destination (dst) where most similar tokens from the src set are merged into the dst set.
-Essentially, tokens in the dst set function as anchors for grouping related tokens and also serve as the primary regions where computation will take place.
-Thus, token partitioning is particularly crucial to minimize performance drop and avoid image artifacts.
-To this end, ToMeSD [3] proposes to select a destination token randomly within each $2\times 2$ region, shown to be better than both fully random sampling and other fixed-pattern methods.
-However, this selection scheme is still random in nature and does not prioritize tokens based on relevance, which potentially misallocates resources to non-essential areas, as can be seen in Fig.
-1 (a).
-Wise old man reading a book by candlelightRefer to caption (a) ToMeSD [3] (b) Ours
+The rise of powerful diffusion models such as DALL-E [79], Stable Diffusion (SD) [83], or Imagen [84] has dramatically changed the landscape of generative AI [27, 94, 83, 84, 2, 18].
+At their core, these models operate by iteratively denoising through multiple passes of a backbone network, processing a substantial number of tokens, which are particularly computationally intensive.
+To reduce the computational demands, prior work [4, 3, 37] has explored merging similar tokens and sharing the results across member tokens.
+Prompt: “A cute owl wearing a wizard hat.Refer to caption (a) Spatial [3](b) Importance-based (Ours)
 Figure 1: Importance-based Token Merging.
@@
 In the second row, we show the regions (in white) where computation (e.g., attention) will take place after token merging.
-To address these limitations, we propose a novel token selection method that focuses computational resources on areas of high relevance to the prompt.
-Unlike the random selection used in ToMeSD [3], our approach leverages classifier-free guidance [25] as a targeted mechanism for token selection.
-While classifier-free guidance has been typically used as a global adjustment to emphasize prompt-relevant features, here we show that we can use it at the token level to dynamically highlight tokens carrying more essential information.
-By prioritizing token selection based on this prompt relevance, we ensure that high-information regions receive focused attention, preserving critical details and enhancing image quality.
-Further, classifier-free guidance incurs no additional computational cost and requires no modification to the architecture, thus, it can be readily used for any conditional diffusion models.
-More specifically, our method assigns each token a relevance score based on its magnitude in the classifier-free guidance map, indicating its importance to the prompt.
-We then construct a pool of important tokens and only select destination tokens from this pool.
-Unlike simply picking the top-k important tokens, which can lead to redundancy and a suboptimal allocation of computing resources, our approach ensures a balance between relevance and diversity, focusing on prompt-aligned areas while maintaining efficiency.
-We apply our token merging strategy across three key applications: text-to-image generation, text-to-video generation, and multi-view image generation.
-In all cases, our method can be integrated seamlessly without requiring any modifications.
-Compared to the baseline method, our approach consistently demonstrates superior performance, delivering results with higher fidelity and significantly improved image details across all tested scenarios.
+However, such a merging process can also significantly degrade image quality, losing important details and structure.
+This typically happens when merging occurs in highly informative image regions, where critical visual details are compressed or discarded.
+One major reason is that existing methods rely on ad hoc heuristics for grouping tokens, often based on spatial proximity or fixed patterns without considering their content.
+Thus, destination tokens, where the other tokens will merge into, are chosen randomly within predefined regions, sometimes pushing important visual elements into less relevant areas.
+This leads to suboptimal resource allocation, where crucial textures, edges, or fine-grained structures are lost, while less important regions retain more detail than necessary, as can be seen in Fig.
+1 (a).
+Existing methods rarely consider token content explicitly when merging.
+To address this limitation, we propose a novel token selection method that prioritizes computational resources in areas of high visual and semantic importance.
+Unlike existing approaches that rely on random or fixed-pattern selection, our method ensures that merging decisions are guided by actual content relevance.
+Rather than treating all tokens equally, we use per-token importance signals to identify key regions and ensure that merging prioritizes preserving more important tokens.
+Note that there are many proxies for importance, such as attention scores, saliency maps, or user-provided bounding boxes, all of which can be integrated into our merging method.
+Specifically in our case, we point out that an excellent choice is classifier-free guidance (CFG) [26], as it inherently highlights regions that strongly influence model outputs and is obtainable with no additional cost.
+More specifically, instead of merging tokens arbitrarily across the image, we first construct a pool of high-importance tokens.
+Then, token partitioning is only performed within this pool using a soft-matching strategy.
+This ensures that the most relevant tokens serve as anchors for merging, preserving key details while also maintaining diversity among the anchor tokens.
+Doing so facilitates a more efficient allocation of computational resources, preventing redundancy and ensuring that merging decisions are both meaningful and effective.
+Compared to simply selecting the top-k important tokens—which can lead to redundancy and suboptimal token assignments—our method strikes a balance between relevance and diversity, resulting in more coherent and detailed generations.
+We apply our token merging strategy across three key applications: text-to-image generation, multi-view generation, and text-to-video generation, as well as across various model structures such as U-Net or transformers.
+Compared to baseline methods, our method consistently demonstrates superior performance, delivering results with higher fidelity and significantly improved image details across all tested scenarios.
 Our contributions are as follows:
-  • We propose using classifier-free guidance magnitude as a measure of token importance. This can be obtained efficiently and robustly without any modification to diffusion models.
-  • We design a novel token-merging strategy based on token importance to improve the quality of generated samples. We introduce an adaptive important token pool to enhance resource allocation.
-  • Our method achieves state-of-the-art performance across various diffusion model tasks, including text-to-image synthesis, multi-view generation, and video generation.
+  • We propose a novel importance-based token merging paradigm for diffusion models, designed to preserve crucial image content.
+The importance score can be easily obtained from CFG at no additional cost.
+  • We design a novel token-partitioning strategy based on a pool of important tokens to improve generation quality.
+  • Our method achieves state-of-the-art performance across various diffusion model tasks, including text-to-image synthesis, multi-view generation, and video generation, as well as across model architectures including U-Net and transformers.
 2 Related Work
@@
 Diffusion models [92, 13, 27, 94] are a class of generative models that iteratively transform random noise into complex data structures, such as images, by gradually reversing a diffusion process.
-In these models, data is progressively corrupted by adding noise, and the model learns to recover the original data through iterative denoising.
-This approach has demonstrated impressive results in generating high-quality, realistic images [78, 79, 70, 6, 74, 66, 16, 64, 104, 1], videos [83, 2, 17, 113, 38, 5], 3D content [72, 51, 56, 82, 41, 59, 67], and audio [36, 27].
+In these models, data is progressively corrupted by adding noise, and the model learns to recover the original data through iterative denoising with learned parameters.
+This approach has demonstrated impressive results in generating high-quality, realistic images [83, 84, 75, 7, 79, 71, 17, 69, 109, 1, 34], videos [88, 2, 18, 118, 40, 6], 3D content [77, 54, 59, 87, 43, 63, 72], and audio [38, 28].
+A popular diffusion model architecture is introduced by Stable Diffusion [83], which uses a U-Net with transformer blocks.
+It first encodes a noisy image as latent tokens, which are processed through transformer blocks comprising self-attention, MLP, and cross-attention layers.
+This design has been extended for multi-view generation with multi-view attention [87] and for video generation with temporal layers [18].
+More recent approaches [7, 118] replace the U-Net with a fully transformer-based architecture for improved scalability.
 In diffusion models, classifier-free guidance (CFG) [26] is a technique that enhances fidelity and detail in generation by guiding the model toward conditioned inputs without the need for an external classifier.
@@
 Some approaches require retraining, including better model structures [83, 76, 36], model pruning [15], model compression [117, 110, 49], step distillation [85, 70, 86, 55, 19], and consistency regularization [95, 64].
-Others avoid retraining, such as improved sampling to reduce inference steps [89, 50, 57, 58], caching [62, 100, 8, 111, 31, 85, 61], model quantization [44, 7, 22, 94, 11, 86], and token reduction [32, 3, 4].
+Others avoid retraining, such as improved sampling to reduce inference steps [94, 53, 60, 61], caching [67, 105, 9, 116, 32, 90, 66, 45], model quantization [47, 8, 23, 99, 12, 91], and token reduction [33, 3, 4].
 In this work, we focus on improving token reduction.
-Notably, our method is compatible with other diffusion acceleration techniques, enabling combined use to further improve inference efficiency.
+Notably, our method is compatible with other diffusion acceleration techniques, enabling combined use for further speedup.
 2.2 Token Reduction
@@
 Additionally, soft merging methods assign tokens to multiple clusters before merging them [20, 82, 119].
-Recent studies have applied token reduction to diffusion model inference [3, 35, 21, 93, 45, 32, 110, 47, 95, 84].
+Recent studies have applied token reduction to diffusion model inference [3, 37, 22, 98, 48, 33, 115, 50, 100, 89, 62].
 ToMeSD [3] introduced token merging to Stable Diffusion [83] with a training-free method.
-It selects destination tokens and uses bipartite soft matching to merge redundant tokens.
-The reduced token set is processed by operations like attention, and then copied back to merged token locations to ensure the final token count matches the input.
-While ToMeSD performs well, its selection of destination tokens is not optimal, i.e., random across the image.
+The typical token merging procedure for diffusion models selects destination tokens from input feature tokens and utilizes bipartite soft matching to merge redundant tokens.
+The reduced token set is processed by operations like attention, and is then copied back to the merged token locations, ensuring the final token count matches the input.
+Although this token merging approach performs well, its selection of destination tokens is not optimal, i.e., spatially random across the image.
+ToFu [37] combines token pruning and merging but still follows similarly suboptimal destination token selection strategy.
 In our work, we propose that selecting destination tokens based on their importance improves the fidelity and quality in generation, and this importance can be easily obtained via classifier-free guidance without additional cost.
 AT-EDM [98] uses self-attention maps to derive token importance, but this requires first performing full self-attention and increases peak memory usage.
-ATC [21] applies bottom-up hierarchical clustering for better token merging, but our experiment shows that this clustering is costly for generative tasks.
-3 Preliminaries
-Diffusion models.
-Diffusion models [12, 26, 89] are generative models that produce data by simulating a denoising process.
-Starting with data $\mathbf{x}_{0}$, the forward diffusion process gradually adds Gaussian noise over $T$ timesteps, producing a sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$.
-This process is defined as:
-$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\,\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
-where $\beta_{t}$ controls the noise schedule.
-The reverse process denoises the input by modeling $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ with learned parameters.
-Most models predict the noise added at each timestep, $\epsilon_{\theta}(\mathbf{x}_{t},t)$.
-Classifier-free guidance (CFG) [25] modifies the noise prediction to improve sample control. It is designed to steer the predicted noise in a direction more aligned with the condition:
-$\tilde{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)),$ (2)
-where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$, $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction, and $w$ is the guidance weight.
-Token Merging for Diffusion Models.
-Our token merging method builds on ToMeSD [3], which is designed to speed up text-to-image diffusion models.
-To accelerate a model layer, such as attention, ToMeSD divides the input feature tokens into a source (src) set and a destination (dst) set.
-The dst set is selected randomly, i.e., by choosing one token from each $2\times 2$ region in the input feature map.
-Using bipartite soft matching, it then selects and merges the $r$ most similar tokens from the src set into the dst set, where $r$ represents the token merging ratio.
-The merging is done via averaging.
-After the model layer processes this reduced token set, the merged locations are filled with the corresponding processed dst tokens, maintaining the same output shape as the input.
+ATC [22] applies bottom-up hierarchical clustering for better token merging, but we show it is costly for generative tasks.
 3 Method
+We propose a novel importance-based token-merging method, summarized in Fig.
+2, allowing for a more efficient allocation of computing resources.
+We highlight that classifier-free guidance serves as an effective importance indicator.
+Further, we propose using a dynamic pool of important tokens, where the pool size adapts to the token merging ratio, optimizing resource allocation and reducing redundancy.
 Figure 2: Overview.
 We propose an importance-based token merging method.
-The importance of each token is determined by the magnitude of classifier-free guidance.
+The importance of each token can be determined using classifier-free guidance.
 These scores, visualized with colors ranging from light to dark (indicating less to more important tokens), are used to construct a pool of important tokens.
@@
 All other src tokens and unimportant tokens are merged with the destination tokens for subsequent computational steps.
-We propose a novel token-merging method for diffusion models.
-Our approach uses classifier-free guidance to identify important tokens, allowing for a more efficient allocation of computing resources.
-Additionally, we create a dynamic pool of important tokens, with the pool size adapting to the token merging ratio.
-This adjustment enhances resource allocation and reduces redundancy when determining which tokens to merge.
-Fig.
-2 provides an overview of our method.
 Important Tokens.
@@
 This is because these tokens correspond to the primary image regions where subsequent computations, such as attention, are applied.
-However, ToMeSD [3] currently relies on a random selection process based on a simple heuristic.
-It lacks well-established criteria for determining token importance.
-The primary guiding principle of ToMeSD is to prioritize spatial uniformity among the destination tokens [3, 4].
+In principle, all reliable per token importance signals, such as cross attention maps, user provided bounding boxes, or saliency maps, can be integrated with our method.
+We find that classifier-free guidance (CFG) [26] serves as an excellent indicator.
+CFG modifies the noise prediction to improve sample control.
+It is designed to steer the predicted noise in a direction more aligned with the condition:
+$\tilde{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)),$ (1)
+where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$, $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction, and $w$ is the guidance weight.
+It is widely used in diffusion models and incurs no additional computational cost.
+For each token, we calculate its importance as the absolute value of its CFG score:
+importance $=\left|\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)\right|$ (2)
+$\approx\left|-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p(y\mid\mathbf{x}_{t})\right|,$
+where $\sigma_{t}$ is the noise scale.
+The guidance term effectively estimates the gradient of the log-likelihood of the conditioning variable $y$ (e.g., a text prompt) with respect to the noisy sample $\mathbf{x}_{t}$ [26].
+Thus, the CFG magnitude can be interpreted as a saliency or importance measure: tokens with a high CFG magnitude have stronger influences in steering the generation toward satisfying the condition $y$.
+In Fig.
+3, we provide an example of the resulting importance maps.
 Fresh citrus and berriesRefer to caption
 Figure 3: Importance Maps.
-We show token importance maps derived from classifier-free guidance across diffusion inference timesteps.
-The generated image from the diffusion model is displayed on the left for reference.
-In this paper, we highlight that classifier-free guidance [25] serves as a reliable indicator of token importance, providing a principled approach to identifying image regions with essential information.
-By leveraging this connection, we propose a method that prioritizes computing resources on important tokens, enabling more efficient and informed token merging even under constrained compute budgets.
-For each token, we calculate its importance as the absolute value of its classifier-free guidance score:
-$\text{importance}=\left|\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)\right|,$ (3)
-where $\epsilon_{\theta}(\mathbf{x}_{t}\mid y,t)$ is the noise prediction conditioned on input $y$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction.
-In Fig.
-3, we provide an example of the importance maps.
-A car made out of vegetablesRefer to caption (a) Top-k (b) Ours
-Figure 4: We compare our method with an approach that uses the top-k important tokens as destination tokens (dst) for token merging.
-The computation locations after token merging are illustrated as non-black pixels in the bottom-left windows.
-They include locations of dst tokens, which are shown in white, and independent tokens (some other tokens that lack a similar dst token for merging), which are shown in orange.
-Figure 5: Qualitative comparison of text-to-image generation.
-The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with various token merging methods including ATC [21], ToMeSD [3], and our proposed method.
-Our approach consistently produces finer details with coherent structures.
-Note that ATC requires minutes to generate an image, whereas other methods, including ours, complete the task in seconds.
-The token merging ratio is 0.7.
-Please refer to the supplementary for detailed prompts.
-Best viewed with zoom-in for clarity.
-Figure 6: Qualitative comparison of multi-view diffusion.
-We apply ToMeSD [3] and our token merging method to the multi-view diffusion model.
-We use Zero123++ [82] as the base model and a merging ratio of 0.6.
-Best viewed with zoom-in.
-Please refer to supplementary for input images.
+We present token importance maps derived from classifier-free guidance (CFG) across diffusion inference timesteps.
+These maps highlight areas significantly align with the user prompt.
+In the early steps, they capture the semantics and structure of the image relevant to the prompt, while in later steps, they focus on finer details of the objects the user intends to generate.
+The generated image is shown on the left for reference.
 Importance-based Token Merging.
-With the token importance scores, a naive approach is to pick the top-k tokens as destination tokens and merge the rest tokens that are similar to them.
+With the token importance scores, a naive approach is to pick the top-k tokens as destination tokens (dst) and merge the rest tokens that are similar to them.
 However, this approach produces low-quality outputs due to merging inefficiency, as shown in Fig.
@@
     Unimportant Independent Tokens.
-In the standard token merging pipeline [3], there are “independent” tokens that don’t have similar destination tokens, which are processed independently without merging.
-In our case, background or irrelevant tokens can easily become independent tokens as there are no good matches for them among the important tokens, as shown in Fig.
+In the token merging pipeline, some tokens lack similar destination tokens, making them “independent” and remain unmerged.
+In the top-k approach, background or irrelevant tokens often become independent due to the absence of suitable matches among the important tokens, as shown in Fig.
 4.
-To this end, we propose to first create a pool of the most important tokens and ensure both destination tokens and independent tokens are drawn from this set.
-Destination tokens are randomly sampled from the pool, while independent tokens are selected as those in the important token pool that are most dissimilar to the chosen destination tokens.
-This simple method ensures that all subsequent computations after the merging step are performed on important tokens, leading to improved detail preservation, as demonstrated in Fig.
+To avoid these issues, we propose to first create a pool of the most important tokens and then, ensure that both destination tokens and independent tokens are drawn from this set.
+To do so, destination tokens are randomly sampled from the pool, while independent tokens are selected as those in the important token pool that are most dissimilar to the chosen destination tokens.
+This approach ensures that all computations following the merging step operate on important tokens, leading to improved detail preservation, as illustrated in Fig.
 4 (b).
@@
 Conversely, with a larger compute budget, the pool size increases, reducing the likelihood of selecting redundant tokens as destination tokens.
+A delicate pink rose in full bloom, detailed petals.Refer to caption (a) Top-k(b) Ours
+Figure 4: We compare our method with an approach that uses the top-k important tokens as destination tokens (dst) for token merging.
+The computation locations after token merging are illustrated as non-black pixels in the bottom-left windows.
+They include locations of dst tokens, which are shown in white, and independent tokens (some other tokens that lack a similar dst token for merging), which are shown in orange.
+Our method produces more structured and detailed image, as highlighted in the red box.
+(a) SD [83] $\sim$ 8s(b) ATC [22] $\sim$ mins(c) ToFu [37] $\sim$ 5s(d) ToMe. [3] $\sim$ 5s(e) Ours $\sim$ 5s
+Figure 5: Qualitative comparison of text-to-image generation.
+The first column shows results from Stable Diffusion (SD) [83], while the subsequent columns show SD combined with various token merging methods.
+As highlighted in red boxes, our approach consistently produces finer details with coherent structures.
+Note that ATC requires minutes to generate an image, whereas other methods, including ours, complete the task in seconds.
+The token merging ratio is 0.7.
+Please see the supplementary for prompts.
+Best viewed with zoom-in.
 Token Merging in Diffusion Inference.
-More specifically, at time-step $t$ of the diffusion inference, a diffusion model layer, e.g.
+At time-step $t$ of the diffusion inference, a diffusion model layer, e.g.
 a transformer layer, takes $\mathbf{N}$ tokens as input.
@@
 From $\mathbf{A}$, we randomly pick $\mathbf{D}=\mathbf{N}\cdot k$ tokens to form the destination (dst) set.
-Here, $p,k$ are the two hyper-parameters of our method.
+Here, $p$, $k$ are hyper-parameters.
 The remaining tokens in $\mathbf{A}$ become the source (src) set.
@@
 The remaining src tokens and unimportant tokens are merged into their corresponding dst tokens.
+The merging is performed via averaging all grouped tokens.
 After merging, the number of tokens reduces to $\mathbf{I}+\mathbf{D}=\mathbf{N}\cdot(1-r)$ tokens.
-Preserving Structure in Early Time-steps.
-Prior work [35] suggests that token pruning (directly dropping tokens) could help preserve structural details.
-Building on this observation, we found that incorporate token pruning during the early diffusion steps, followed by token merging in later steps, improves generation results.
-Furthermore, the low token variance in the early steps [93] reduces the effectiveness of classifier-free guidance in identifying important tokens.
-To mitigate this, we simply employ the ToMeSD strategy during these initial steps, e.g., the first 6 out of 50 steps, to select one token from each $2\times 2$ region of the feature map as a destination token.
+The diffusion model layer then processes this reduced token set, the merged locations are filled with the corresponding processed dst tokens, maintaining the same output shape as the input.
 4 Experiments
-Our token merging strategy leverages classifier-free guidance for relevance-based token selection, ensuring an efficient and seamless integration without requiring any additional modules or modifications to diffusion models.
-To showcase its effectiveness, we evaluate our method across three diffusion generation tasks: text-to-image generation, image-conditioned multi-view diffusion, and text-to-video diffusion.
 4.1 Experimental Settings
@@
 We use Stable Diffusion 2 (SD) [83] as the base model, a text-to-image latent diffusion model with a U-Net architecture that generates 768x768 images.
-Our token merging method is compared to ToMeSD [3] and ATC [21].
+Our token merging method is compared to ToFu [37], ToMeSD [3] and ATC [22].
 Due to ATC’s slow inference (several minutes per image), we only display its visual results.
 For quantitative comparison, we follow previous studies [84, 109, 1, 79] and report FID [25] and CLIP scores [78, 24] for zero-shot image generation on the MS-COCO 2014 validation dataset [52], with 30K randomly sampled image-caption pairs.
+We also experiment with a diffusion transformer, PixArt-$\alpha$ [7], for text-to-image generation.
+We compare our method with ToMeSD and similarly evaluate on the MS-COCO dataset.
+Unless otherwise stated, “text-to-image” in this paper refers to generation based on Stable Diffusion.
 Multi-view Diffusion.
@@
 Zero123++ is an image-conditioned multi-view latent diffusion model that generates six novel views at a resolution of 320×320.
-We compare our token merging method to ToMeSD.
-Following the evaluation protocols of prior work [53, 56], we test on GSO dataset [13], which comprises 30 everyday objects, and compute PSNR, SSIM [98], and LPIPS metrics [108] to evaluate the similarity between the generated images and ground truth.
-Additionally, we include visual comparisons using in-the-wild images as input to further demonstrate the effectiveness of our method.
+We compare our method to ToMeSD.
+Following evaluation protocols of prior work [56, 59], we test on GSO dataset [14], which comprises 30 everyday objects, and compute PSNR, SSIM [103], and LPIPS metrics [113] to evaluate the similarity between the generated images and ground truth.
+We also include visual comparisons using in-the-wild images as input.
+(a) PixArt-$\alpha$ [7](b) ToMeSD [3](c) Ours
+Figure 6: Token merging for diffusion transformer.
+We apply ToMeSD [3] and our method to PixArt-$\alpha$ [7] with a merging ratio of 0.3.
+Detailed generations are highlighted with red boxes.
+Best viewed with zoom-in.
+Please see the supplementary for prompts.
+Figure 7: Qualitative comparison of multi-view diffusion.
+We apply ToMeSD [3] and our token merging method to the multi-view diffusion model.
+We use Zero123++ [87] as the base model and a merging ratio of 0.6.
+Our method outputs finer details, as highlighted in red boxes.
+Best viewed with zoom-in.
+Please refer to the supplementary for input images.
 Video Diffusion.
@@
 Other Metrics.
-We also report TFLOPs, latency, and GPU memory usage, with and without memory-efficient attention (xFormers [40]), on an NVIDIA A5000 GPU.
-Inference is performed using float16 precision.
+We report TFLOPs, latency, and GPU memory usage, with and without memory-efficient attention [42], on an NVIDIA A5000 GPU.
+Inference uses float16 precision.
 We estimate TFLOPs for a single diffusion sampling step.
@@
 Implementation Details.
-For text-to-image diffusion, we follow the approach used in ToMeSD [3] by merging tokens in the self-attention layers of the first and last blocks of Stable Diffusion.
-The diffusion process consists of 50 sampling steps, with token pruning applied during the first 6 steps, followed by token merging.
-The classifier-free guidance (CFG) scale is set to 7.5.
+For text-to-image synthesis with Stable Diffusion [83], we merge tokens in the self-attention layers of the first and last model blocks, similar to ToMeSD [3].
+For PixArt-$\alpha$ [7], we apply token merging to self-attention and cross-attention in the middle half of the model layers (7–20 out of 28) for simplicity.
 For multi-view diffusion, we merge tokens in the self-attention layers of the first two and last two blocks of Zero123++ [87].
-The process involves 50 sampling steps, with token pruning applied in the first 10 steps.
-The classifier-free guidance (CFG) scale is set to 4.
-For video diffusion, token merging is applied in the first and last blocks of AnimateDiff [17].
-The sampling consists of 30 steps, with token pruning during the first 4 steps and a CFG scale of 7.5.
-To ensure fairness, these settings are consistently applied to both our method and the baselines.
-The hyper-parameter $p$, which determines the size of the important token pool, is set to 0.4, 0.6, and 0.8 for image, multi-view, and video generation task, respectively.
-The number of destination tokens ($k$) is consistent across all tasks and follows the setting used in ToMeSD, utilizing 25% of the total tokens.
+For video diffusion, we merge tokens in the first and last blocks of AnimateDiff [18].
+The hyper-parameter $p$, which determines the size of our important token pool, is set to 0.4, 0.6, and 0.8 for image, multi-view, and video generation tasks, respectively.
+The number of destination tokens ($k$) remains consistent across all tasks and follows the setting used in ToMeSD, utilizing 25% of the total tokens.
 4.2 Results
-Figure 7: Qualitative comparison of video diffusion.
-We apply ToMeSD and our token merging method to the video diffusion model.
-For each generated video, we show three frames: the first on the left, the 8th at the top right, and the last 16th frame at the bottom right.
-We use AnimateDiff [17] as the base model and a merging ratio of 0.7.
-Best viewed with zoom-in.
-Please refer to supplementary for prompts.
+In Tab.
+1 and Fig.
+5, we compare different token merging methods applied to Stable Diffusion 2 [83] for text-to-image generation.
+Our method consistently outperforms baselines, especially at higher token merging ratios ($r$).
+For example, at $r=0.75$, our method achieves an FID of 17.75 versus ToMeSD [3]’s 20.89.
+Agglomerative Token Clustering (ATC) [22] is not included due to its prohibitive computational cost, as it is CPU-bound and non-batched, making it impractical for large-scale evaluations.
+Notably, our method also performs well when used with cross-attention maps (Sec.
+4.3).
+When the merging ratio is small, such as 0.3, our important token pool becomes the whole token set, making it functionally equivalent to ToMeSD.
 $r$ FID $\downarrow$ CLIP $\uparrow$
-ToMeSD Ours ToMeSD Ours
-0 11.88 31.83
-0.30 12.20 12.20 31.82 31.82
-0.50 13.50 13.42 31.79 31.83
-0.60 14.81 14.51 31.80 31.81
-0.70 17.46 16.22 31.78 31.79
-0.75 20.89 17.75 31.71 31.76
-Table 1: Text-to-image generation. We apply ToMeSD [3] and our token merging method to Stable Diffusion [78] across various token merging ratios $r$.
-In Tab.
-1, we compare different token merging methods applied to Stable Diffusion 2 [78] for text-to-image generation.
-Our method consistently outperforms ToMeSD [3] in this task, especially at higher token merging ratios ($r$).
-For example, at $r=0.75$, our method achieves an FID of 17.75 versus ToMeSD’s 20.89.
-Agglomerative Token Clustering (ATC) [21] is not included due to its prohibitive computational cost, as it is CPU-bound and non-batched, making it impractical for large-scale evaluations.
-We provide qualitative comparisons in Fig.
-5.
-Note that when the token merging ratio $r$ is small, such as 0.3, our important token pool becomes the whole token set, making it functionally equivalent to ToMeSD.
+ToFu ToMe. Ours ToFu ToMe. Ours
+0 - - 11.88 - - 31.83
+0.30 13.48 12.20 12.20 31.82 31.82 31.82
+0.50 15.81 13.50 13.42 31.81 31.79 31.83
+0.60 17.32 14.81 14.51 31.81 31.80 31.81
+0.70 18.94 17.46 16.22 31.78 31.78 31.79
+0.75 19.29 20.89 17.75 31.76 31.71 31.76
+Table 1: Text-to-image generation. We apply ToFu [37], ToMeSD [3] and our token merging method to Stable Diffusion [83] across various token merging ratios $r$.
+$r$ FID $\downarrow$ CLIP $\uparrow$ Latency (s) $\downarrow$
+ToMe. Ours ToMe. Ours
+0 27.51 31.30 9.14
+0.3 34.23 28.50 30.76 31.05 8.96
+0.5 65.46 34.02 29.85 30.68 7.54
+0.7 95.46 50.76 28.67 29.93 7.14
+Table 2: Comparison of our method with ToMeSD [3] when applied to the diffusion transformer PixArt-$\alpha$ [7].
 $r$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
 0.40 14.80 14.82 0.775 0.777 0.260 0.259
-0.60 14.71 14.83 0.782 0.783 0.272 0.269
+0.60 14.71 14.85 0.782 0.783 0.272 0.263
 0.70 14.18 14.80 0.787 0.785 0.302 0.274
 0.75 13.12 14.58 0.789 0.784 0.349 0.283
-Table 2: Multi-view diffusion. We apply ToMeSD [3] and our token merging method to Zero123++ [82] with different token merging ratios $r$.
+Table 3: Multi-view diffusion. We compare ToMeSD [3] with our token merging method when applied to Zero123++ [87].
 In Tab.
-2, we compare ToMeSD and our token merging method in the context of multi-view diffusion.
+2 and Fig.
+6, we show our method significantly outperforms ToMeSD when applied to a diffusion transformer, achieving an FID improvement of 17–48%.
+This highlights the generalizability of our approach.
+In Tab.
+3, we compare ToMeSD and our token merging method in the context of multi-view diffusion.
 Our importance-based token merging method consistently shows improved performance over ToMeSD, especially at higher merging ratios.
@@
 7 further validate this, showcasing finer geometrical and textual details in the objects generated by our method.
+A similar trend can be observed in Tab.
+4 and Fig.
+8, where we extend the comparison to video diffusion.
+Across these tests, our method consistently performs better than ToMeSD, both in numerical metrics and in visual quality, particularly in preserving object details.
+We observed that merging tokens in the temporal layers significantly reduces the generated dynamics.
+Nonetheless, we present results with token merging applied for both spatial and temporal layers in Tab.
+5.
 $r$ Semantic $\uparrow$ Quality $\uparrow$ Total $\uparrow$
@@
 We report VBench scores [29] of our method in comparison to ToMeSD [3].
-A similar trend can be observed in Tab.
-3 and Fig.
-7, where we extend the comparison to video diffusion.
-Across these tests, our method consistently performs better than ToMeSD, both in numerical metrics and in visual quality, particularly in preserving object details.
-We observed that merging tokens in the temporal layers significantly reduces the generated dynamics.
-Nonetheless, we present results with token merging applied for both spatial and temporal layers in Tab.
-4.
-In summary, our method outperforms ToMeSD across all three tasks, generating images and videos with better object details.
+Figure 8: Qualitative comparison of video diffusion.
+We apply ToMeSD and our token merging method to the video diffusion model.
+For each generated video, we show three frames: the first on the left, the 8th at the top right, and the last 16th frame at the bottom right.
+We use AnimateDiff [18] as the base model and a merging ratio of 0.7.
+Best viewed with zoom-in.
+Please refer to supplementary for prompts.
+Inference Costs.
+We compare the inference costs of our method and ToMeSD when applied to Stable Diffusion 2 in Tab.
+6.
+As can be seen, both methods show similar improvements in inference times, GPU usage, and TFLOPs.
+This demonstrates that our token merging strategy can enhance performance without incurring additional costs.
 $r$ Latency (s) Memory (GB) TFLOPs
@@
 Numbers in parentheses indicate measurements when memory-efficient attention [42] is enabled.
-Inference Costs.
-We compare the inference costs of our method and ToMeSD when applied to Stable Diffusion 2 in Tab.
-5.
-As can be seen, both methods show similar improvements in inference times, GPU usage, and TFLOPs.
-This demonstrates that our token merging strategy can enhance performance without incurring additional computational costs.
 4.3 Ablation Study
+We conduct ablation studies on the text-to-image generation task to examine alternative design choices in our method.
+As shown in Tab.
+7, simply using top-k important tokens as destination tokens leads to worse results.
+Furthermore, not choosing independent tokens exclusively from the important set (w/ global ind.), leads to a performance drop.
 $r$ Ours w/ top-k dst w/ global ind.
@@
 CLIP $\uparrow$ 31.82 31.82 31.82
-0.6 FID $\downarrow$ 14.51 14.56 14.73
-CLIP $\uparrow$ 31.81 31.83 31.82
-Table 6: Ablation studies on alternative design choices of our method for text-to-image generation.
+0.7 FID $\downarrow$ 16.22 16.29 16.43
+CLIP $\uparrow$ 31.79 31.79 31.80
+Table 7: Ablation studies of our method for text-to-image generation.
 ’w/ top-k dst’ means top-k rather than random selection from the important token pool for destination tokens.
 ’w/ global ind.’ means independent tokens may also be outside the important token pool, instead of solely within it.
+$r$ FID $\downarrow$ CLIP $\uparrow$
+ToMe Ours Ours ToMe Ours Ours
+(CA) (CFG) (CA) (CFG)
+0.30 12.20 12.17 12.20 31.82 31.82 31.82
+0.50 13.50 13.38 13.42 31.79 31.82 31.83
+0.70 17.46 16.79 16.22 31.78 31.79 31.79
+0.75 20.89 18.17 17.75 31.71 31.75 31.76
+Table 8: Comparison of token merging methods applied to Stable Diffusion [83]. CA and CFG denote cross-attention and classifier-free guidance as importance signals, respectively.
 $r$ $p$ 0 0.2 0.4 0.8
@@
 Table 9: Choice of $p$. We show the results of our method for the text-to-image generation task with different values of $p$, which determines the important token pool size.
-We conduct ablation studies on the text-to-image generation task to examine alternative design choices in our method.
+Cross-Attention Maps for Token Importance.
+In principle, our method can be used with any per-token importance scores.
 As shown in Tab.
-6, simply using top-k important tokens as destination tokens leads to worse results.
-Furthermore, not choosing independent tokens explicitly from the important set (w/ global ind.), leads to a performance drop.
+8, we use our method with cross-attention maps instead of CFG.
+While this approach may require more memory compared to CFG, it remains a strong alternative and highlights the generalizability of our method.
 Choice of $p$.
-As described in Sec.
-4, the size of the important token pool is calculated as $(1+r)\cdot(1+p)$, where $r$ represents the token merging ratio.
-In Tab.
-7, we present results for text-to-image generation using different values of $p$.
-Setting $p$ too low results in an overly small pool of important tokens, which increases redundancy when randomly selecting destination tokens from it.
-Conversely, setting $p$ too high includes too many unimportant tokens in the pool, negatively impacting performance.
+Our important token pool size is $(1-r)\cdot(1+p)$, where $r$ is the token merging ratio.
+Tab.
+9 shows that our method remains robust to the choice of $p$, provided that $p$ is within a reasonable range.
 5 Conclusions
-In this paper, we propose an importance-based token merging method for diffusion models, which maintains generation quality while reducing inference latency.
-Leveraging classifier-free guidance as an indicator of token importance, our approach strategically allocates computational resources to regions of high relevance to the input condition, thereby enhancing the fidelity of the generated outputs.
+We propose an importance-based token merging method for generation tasks, which maintains generation quality while reducing inference latency.
+We utilize token importance to strategically allocate computational resources to regions of high relevance to the input condition, thereby enhancing the fidelity of the generated outputs.
+This novel, simple, and intuitive strategy accelerates various models for free with no modifications needed.
+Notably, we identify classifier-free guidance as an effective token importance indicator.
 Our method achieves state-of-the-art performance across diverse tasks, including text-to-image synthesis, multi-view generation, and video generation, highlighting its effectiveness and versatility.
-Our approach reveals the surprising effect of applying classifier-free guidance at the token level to facilitate compression with minimal impact on quality.
 Supplementary Material
-In Appendix A, we present more qualitative comparisons, results on preserving output structure, and experiments with varying diffusion inference steps.
+In Appendix A, we present more qualitative comparisons, empirical evidence on the relationship between CFG and token importance, results on consistency models, results on combining our method with orthogonal diffusion acceleration techniques, and experiments with varying diffusion inference steps.
 In Appendix B, we provide more details about our experimental settings.
@@
 Additional Qualitative Results.
-We provide additional qualitative comparisons between our token merging method and ToMeSD [3].
-Additional visual comparisons for text-to-image generation are shown in Fig.
-8, and results for multi-view diffusion are presented in Fig.
-9.
+In Fig.
+9, we provide additional qualitative comparisons between ToFu [37], ToMeSD [3], our token merging method, and the variant of our method using cross-attention maps as importance signals.
+Additional visual comparisons for token merging applied to diffusion transformer are shown in  Fig.
+10.
+Additional results for multi-view diffusion are presented in  Fig.
+11.
 In the supplementary video, we include comparisons on text-to-video generation, using AnimateDiff [18] as the base diffusion model and a merging ratio of 0.7.
 Furthermore, we provide visual comparisons between ToMeSD [3] and our method across various merging ratios for text-to-image generation in Fig.
-10.
-Preserving Structure in Early Timesteps.
+12.
+Token Importance via Classifier-Free Guidance.
+The absolute value of classifier-free guidance (CFG) can be interpreted as a token-level saliency measure, highlighting the tokens that play a crucial role in steering the output toward the given prompt or condition.
+Empirically, as shown in Tab.
+10, removing the top 30% of high-CFG tokens leads to a significant degradation in generation results, whereas removing the bottom 30% has little impact.
+No pruning Top 30% Bottom 30%
+FID $\downarrow$ 11.88 15.48 12.78
+CLIP $\uparrow$ 31.83 31.58 31.88
+Table 10: Comparison of pruning (dropping) the top 30% of tokens with the highest CFG values versus the bottom 30% during text-to-image generation using Stable Diffusion [83].
+Results on the Consistency Model.
+Consistency models [95, 64, 65] typically distill classifier-free guidance (CFG), making the explicit guidance term inaccessible, as they approximate the final guided noise function within a single forward pass.
+However, our method is not limited to CFG and can leverage any reliable per-token importance signal.
+As shown in Tab.
+11, our token merging, using cross-attention maps as importance signals, remains effective and outperforms the baseline model when applied to the latent consistency model [65].
+$r$ FID $\downarrow$ CLIP $\uparrow$ Time Mem.
+ToMe. Ours ToMe. Ours (s) $\downarrow$ (GB)$\downarrow$
+0 25.38 31.05 0.65 6.75
+0.30 25.63 25.63 31.03 31.03 0.60 4.22
+0.50 27.80 27.51 30.98 30.99 0.51 3.56
+0.60 30.05 29.61 30.90 30.92 0.50 3.37
+0.70 35.01 32.49 30.59 30.77 0.48 3.21
+0.75 43.28 36.66 30.38 30.60 0.46 3.12
+Table 11: Results on the Consistency Model. We show the comparison of token merging methods applied to a 4-step latent consistency model (LCM_Dreamshaper_v7) [65] across various merging ratios $r$.
+Combination with Orthogonal Diffusion Acceleration Methods.
+In Tab. 12, we demonstrate that our method can be combined with orthogonal diffusion acceleration methods [67, 45, 90] to further accelerate inference while preserving generation quality.
+FID $\downarrow$ CLIP $\uparrow$ Time Mem.
+(s) $\downarrow$ (GB) $\downarrow$
+Ours 16.22 31.79 5.8 3.55
++ DeepCache [67] 15.46 31.81 2.6 3.56
++ FasterDiff. [45] 12.48 31.80 4.2 6.33
++ FRDiff [90] 15.25 31.83 3.5 3.59
+Table 12: Combination with Diffusion Acceleration Methods.
+We show text-to-image generation results by integrating our method with orthogonal diffusion acceleration techniques, using Stable Diffusion [83] as the base model and a merging ratio of 0.7.
+Number of Diffusion Inference Steps
+In Tab. 13, we compare ToMeSD [3] and our method across different numbers of diffusion inference steps for text-to-image generation, demonstrating that our method consistently outperforms ToMeSD.
+$T$ FID $\downarrow$ CLIP $\uparrow$
+ToMeSD Ours ToMeSD Ours
+20 18.57 17.03 31.77 31.80
+30 17.82 16.51 31.78 31.80
+50 17.46 16.22 31.78 31.79
+Table 13: Number of Diffusion Inference Steps.
+We compare ToMeSD [3] and our token merging method for the text-to-image generation task with different diffusion inference steps, using Stable Diffusion [83] as the base model and a merging ratio of 0.7.
+Appendix B Additional Experimental Settings
+Metrics.
+We use the deepspeed [81] library to estimate TFLOPs, the clean-fid [74] library to calculate FID scores, and the openai/clip-vit-base-patch16 model from OpenAI-CLIP [78] to calculate CLIP scores.
+Additional Implementation Details.
+For text-to-image generation using Stable Diffusion [83], the diffusion process consists of 50 sampling steps, with the CFG scale set to 7.5.
+For PixArt-$\alpha$ [7], we perform diffusion sampling for 20 steps, with a CFG scale of 4.5.
+When computing token similarity for token merging in PixArt-$\alpha$, we find that using pixel location distance of tokens yields better results than feature similarity, and we adopt this approach.
+For multi-view diffusion using Zero123++ [87], the process involves 50 sampling steps, with the CFG scale set to 4.
+For video diffusion using AnimateDiff [18], the sampling consists of 30 steps, with a CFG scale of 7.5.
+To ensure fairness, these settings are consistently applied to both our method and the baselines.
+For the ablation study investigating the use of cross-attention maps as importance signals, we utilize the averaged attention map from the final model block in Stable Diffusion.
+Details on Multi-view Diffusion.
+The base multi-view diffusion model used in our experiments is Zero123++ [87], which fine-tunes Stable Diffusion 2 [83] to generate six novel views from an input image.
+During denoising, the model appends the self-attention key and value matrices from the reference input image to the attention layers for conditioning.
+The novel view poses are defined by a fixed set of absolute elevation and relative azimuth angles.
+Specifically, the elevation and azimuth angles (in degrees) are set as follows: (30, 30), (-20, 90), (30, 150), (-20, 210), (30, 270), (-20, 330).
+We consistently use this sequence of novel views to present visual results in multi-view diffusion experiments.
+Preserving Structure in Early Time-steps.
+Prior work [37] suggests that token pruning (directly dropping tokens) could help preserve structural details.
+Building on this observation, we found that incorporate token pruning during the early diffusion steps, followed by token merging in later steps, improves generation results.
+Specifically, we apply token pruning during the first 6, 10, and 4 diffusion inference steps for image, multi-view, and video generation, respectively.
+Furthermore, the low token variance in the early steps [98] reduces the effectiveness of classifier-free guidance in identifying important tokens.
+To mitigate this, during these initial steps that involve token pruning, we randomly select one token from each $2\times 2$ region of the feature map as the destination token.
 In Tab.
 14 and Fig.
-11, we compare the results of ToMeSD [3] under two scenarios: (1) applying token pruning during the early diffusion steps followed by token merging, and (2) using token merging throughout all diffusion steps.
+13, we compare the results of ToMeSD [3] under two scenarios: (1) applying token pruning during the early diffusion steps followed by token merging, and (2) using token merging throughout all diffusion steps.
 The results show that token pruning in the early steps more effectively preserves the generation layout.
+In Tab.
+15, we show that pruning tokens during the first 5-20% denoising steps consistently improves performance across different tasks, highlighting its robustness.
 $r$ FID $\downarrow$ CLIP $\uparrow$
@@
 We evaluate using the text-to-image generation task with Stable Diffusion [83] as the base model across various token merging ratios $r$.
-Number of Diffusion Inference Steps
-In Tab. 9, we compare ToMeSD [3] and our method across different numbers of diffusion inference steps for text-to-image generation, demonstrating that our method consistently outperforms ToMeSD.
-$T$ FID $\downarrow$ CLIP $\uparrow$
-ToMeSD Ours ToMeSD Ours
-20 18.57 17.03 31.77 31.80
-30 17.82 16.51 31.78 31.80
-50 17.46 16.22 31.78 31.79
-Table 9: Number of Diffusion Inference Steps.
-We compare ToMeSD [3] and our token merging method for the text-to-image generation task with different diffusion inference steps, using Stable Diffusion [78] as the base model and a merging ratio of 0.7.
+Metric Pruning Steps (m%)
+0% 5% 10% 20%
+Image
+   FID $\downarrow$ 16.27 16.28 16.22 16.13
+   CLIP $\uparrow$ 31.77 31.79 31.79 31.79
+Multi-View
+   PSNR $\uparrow$ 14.73 14.95 14.75 14.80
+   SSIM $\uparrow$ 0.783 0.782 0.787 0.785
+   LPIPS $\downarrow$ 0.279 0.269 0.268 0.274
+Video
+   Score $\uparrow$ 79.36 79.59 79.63 79.66
+Table 15: We apply token pruning to the first $m\%$ denoising steps (merging ratio = 0.7) and evaluate on three generation tasks.
+Appendix C Discussion and Limitations
+Our method demonstrates broad applicability across diffusion models.
+For multi-guidance scenarios (e.g., InstructPix2Pix [5]), weighted averaging of importance signals based on user preferences could be beneficial.
+For step-distilled diffusion models [70], which already distills classifier-free guidance into the final model, our approach can still be applied by utilizing alternative importance signals, such as attention maps.
+However, an interesting direction for future research could involve refining the distillation process to enable the model to predict an additional output: a classifier-free guidance map, which could then be used for better token merging or other innovative applications.
+Appendix D Prompts
+We provide the prompts used to generate the qualitative results shown in the paper but not included in the figures.
+Text prompts corresponding to the text-to-image generations in Figure 5 of the main paper:
+  • Elegant teacup with a delicate floral pattern
+  • Young musician playing guitar on stage
+  • Colorful butterfly with wings fully spread
+Text prompts corresponding to the text-to-image generations in Figure 6 of the main paper:
+  • A cute cat
+  • A real beautiful face
+  • A small cactus with a happy face in the Sahara desert
+Text prompts corresponding to the text-to-video generations in Figure 8 of the main paper:
+  • Tower
+  • A jellyfish floating through the ocean, with bioluminescent tentacles
+  • In a still frame, the ornate Victorian streetlamp stands solemnly, adorned with intricate ironwork and stained glass panels
+In Fig. 14, we show image prompts corresponding to the image-conditioned multi-view generations in Figure 7 of the main paper.
+In Fig. 15, we show image prompts corresponding to the image-conditioned multi-view generations in Fig. 11.
 Figure 9: Additional comparison of text-to-image generation.
-The first row shows results from Stable Diffusion (SD) [78], while the subsequent rows show SD combined with ToMeSD [3] and our token merging method.
+The first row shows results from Stable Diffusion (SD) [83], while the subsequent rows show SD combined with ToFu [37], ToMeSD [3], our method using cross-attention (CA) map, and our method using classifier-free guidance.
 The token merging ratio is 0.7.
+Our method outputs finer details, as highlighted in red boxes.
+Notably, the variant of our method that utilizes the cross-attention map also achieves better generation details compared to baseline methods, demonstrating the generalization ability of our method.
+Best viewed with zoom-in for clarity.
+(a) PixArt-$\alpha$(b) ToMeSD(c) Ours
+Figure 10: Additional comparison for token merging applied to diffusion transformer.
+We apply ToMeSD [3] and our token merging method to PixArt-$\alpha$ [7] for text-to-image synthesis, using a token merging ratio of 0.4.
+We highlight our generation details with red boxes.
 Best viewed with zoom-in for clarity.
 Figure 11: Additional qualitative comparison of multi-view diffusion.
-We apply ToMeSD [3] and our token merging method to the multi-view diffusion model.
-We use Zero123++ [82] as the base model and a merging ratio of 0.6.
+Token merging is applied to the multi-view diffusion model, Zero123++ [87], with merging ratio as 0.6.
+Our method outputs finer details, as highlighted in red boxes.
 Best viewed with zoom-in.
@@
 For reference, the results of Stable Diffusion without token merging are shown on the left.
+Our method outputs finer details, as highlighted in red boxes.
 Best viewed with zoom-in for clarity.
@@
 We use Stable Diffusion [83] as the base model and a merging ratio of 0.6.
-Appendix B Additional Experimental Settings
-Metrics.
-We use the deepspeed [76] library to estimate TFLOPs, the clean-fid [69] library to calculate FID scores, and the openai/clip-vit-base-patch16 model from OpenAI-CLIP [73] to calculate CLIP scores.
-Details on Multi-view Diffusion.
-The base multi-view diffusion model used in our experiments is Zero123++ [82], which fine-tunes Stable Diffusion 2 [78] to generate six novel views from an input image.
-During denoising, the model appends the self-attention key and value matrices from the reference input image to the attention layers for conditioning.
-The novel view poses are defined by a fixed set of absolute elevation and relative azimuth angles.
-Specifically, the elevation and azimuth angles (in degrees) are set as follows: (30, 30), (-20, 90), (30, 150), (-20, 210), (30, 270), (-20, 330).
-We consistently use this sequence of novel views to present visual results in multi-view diffusion experiments.
-Appendix C Discussion and Limitations
-Our method demonstrates broad applicability across diffusion models.
-However, step-distilled diffusion models [65] do not explicitly compute classifier-free guidance, as this information is already distilled into the final model.
-An interesting direction for future research could involve refining the distillation process to enable the model to predict an additional output: a classifier-free guidance map, which could then be used for better token merging or other innovative applications.
-Appendix D Prompts
-We provide the prompts used to generate the qualitative results shown in the paper but not included in the figures.
-Text prompts corresponding to the text-to-image generations in Fig. 5:
-  • Colorful butterfly with wings fully spread
-  • Elegant teacup with a delicate floral pattern
-  • Royal castle with golden towers at sunrise
-  • Young musician playing guitar on stage
-Text prompts corresponding to the text-to-video generations in Fig. 7:
-  • Tower
-  • A jellyfish floating through the ocean, with bioluminescent tentacles
-  • In a still frame, the ornate Victorian streetlamp stands solemnly, adorned with intricate ironwork and stained glass panels
-In Fig. 12, we show image prompts corresponding to the image-conditioned multi-view generations in Fig. 6.
+Figure 14: Image prompts for Figure 7 of the main paper.
 Figure 15: Image prompts for Fig. 11.
-In Fig. 13, we show image prompts corresponding to the image-conditioned multi-view generations in Fig. 9.
-Figure 13: Image prompts for Fig. 9.